OpenStudy (anonymous):

What is the maximum number of electrons that can be contained in the first, second, third, and fourth energy levels, respectively? Express your answer as integers separated by commas.

OpenStudy (anonymous):

2n^2 electrons in each energy level n is for the shell number

OpenStudy (anonymous):

How would you write that? It wants me to separate with commas?

OpenStudy (anonymous):

if u want to find elections in 3dr shell, n=3 2(3)^2=2*9=18

OpenStudy (anonymous):

but in last she there shud b 8 or less electrons

OpenStudy (anonymous):

2,8,18,32,50,.....